OpenStudy (anonymous):

On a certain committee there are 7 senators. Three of these members will be chosen to form a subcommittee. How many possible subcommittees are there?

OpenStudy (anonymous):

you are being asked for "7 choose 3" or \(\binom{7}{3}\) do you know how to compute it ?

OpenStudy (anonymous):

No, and we can't use any calcualtors

OpenStudy (anonymous):

it is not hard make a fraction on the top put 3 whole numbers starting and 7 and counting down in the bottom, put \(3!=3\times 2\)

OpenStudy (anonymous):

you get \[\frac{7\times 6\times 5}{3\times 2}\] cancel first, multiply last

OpenStudy (anonymous):

I understand the concept behind the numerator but i don't understand why the denominator is what it is...

OpenStudy (anonymous):

because if you just compute \[7\times 6\times 5\] that counts choosing senators A, B, C different from choosing A, C, B B, A, C B, C, A C, A, B C, B, A i.e. it counts each selection 6 times, the number of ways you can arrange 3 senators so you have to divide by 6

OpenStudy (anonymous):

similarly for example \[\binom{10}{4}=\frac{10\times 9\times 8\times 7}{4\times 3\times 2}\]

OpenStudy (anonymous):

Ok so the way to find the denominator for this type of problem you just multiply the number your choosing by each descending whole number down to 1?

OpenStudy (anonymous):

yes, that is called "factorial"

OpenStudy (anonymous):

ok, thank you so much!

OpenStudy (anonymous):

don't forget that the answer to "how many?" is always a whole number, so cancel first multiply last there will never be a fraction when you are done, onlyi a whole number

OpenStudy (anonymous):

yw